OpenStudy (jaylelile):

Which equation represents an exponential function that passes through the point (2, 36)? f(x) = 4(3)x f(x) = 4(x)3 f(x) = 6(3)x f(x) = 6(x)3

OpenStudy (jaylelile):

@nincompoop ?

OpenStudy (loser66):

Any idea?

OpenStudy (jaylelile):

not at all...

OpenStudy (loser66):

Exponential function means the variable is on the exponent of the expression. Can you see which of the options have x on the exponent?

OpenStudy (jaylelile):

the first and 3rd one

OpenStudy (loser66):

good, Now, with the point (2,36), you have if x =2, f(x) =36 between option 1 and 3, which satisfies that condition? (just replace x =2 into it and calculate the result)

OpenStudy (jaylelile):

it's the 3rd one! right?

OpenStudy (loser66):

Please, don't guess!!! show me your work.

OpenStudy (jaylelile):

no no I did calculate. I did 6x3x2=36. did I not do it right?

OpenStudy (loser66):

\(3^2=3*3\) it is not 3*2

OpenStudy (jaylelile):

ohhhhhhhh okay I understand now. Thank you very much

OpenStudy (loser66):

ok. So, which one is the final result?

OpenStudy (jaylelile):

f(x) = 4(3)^x

OpenStudy (loser66):

yup